OpenStudy (anonymous):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = 1/3 x + 7 y = -3x + 3 y = 1/3 x + 3 y = -3x + 7

OpenStudy (solomonzelman):

the slope will remain the same, but the y-intercept will change. let this new y-intercept be represented using "b" so, y=1/3x+b we know it (the new line) passes through (-1,6) so plug it in (to find the new y-intercept e.i. "b" ) (6)=(1/3) (-1) + b 6= - 1/3 +b 6 1/3 = b knowing the new y-intercept the new equation is, y=1/3x+ 6 1/3

OpenStudy (anonymous):

So C!!!!!

OpenStudy (solomonzelman):

with the conditions you posted, it's y = 1/3x+ 6 1/3 and nothing else.

OpenStudy (solomonzelman):

maybe it's supposed to be perpendicular, or the point you entered is entered incorrectly ?